# Direct Superpoints Matching for Fast and Robust Point Cloud Registration

### v1

@@ Abstract
Although deep neural networks endow the downsampled superpoints with discriminative feature representations, directly matching them is usually not used alone in state-of-the-art methods, mainly for two reasons. First, the correspondences are inevitably noisy, so RANSAC-like refinement is usually adopted. Such ad hoc postprocessing, however, is slow and not differentiable, which can not be jointly optimized with feature learning. Second, superpoints are sparse and thus more RANSAC iterations are needed. Existing approaches use the coarse-to-fine strategy to propagate the superpoints correspondences to the point level, which are not discriminative enough and further necessitates the postprocessing refinement. In this paper, we present a simple yet effective approach to extract correspondences by directly matching superpoints using a global softmax layer in an end-to-end manner, which are used to determine the rigid transformation between the source and target point cloud. Compared with methods that directly predict corresponding points, by leveraging the rich information from the superpoints matchings, we can obtain more accurate estimation of the transformation and effectively filter out outliers without any postprocessing refinement. As a result, our approach is not only fast, but also achieves state-of-the-art results on the challenging ModelNet and 3DMatch benchmarks. Our code and model weights will be publicly released.
1 Introduction
@@
Point cloud registration refers to the task of aligning two partially overlapping point clouds into a shared coordinate system. In this paper, we tackle the problem of rigid registration where the goal is to determine the optimal transformation matrix, including rotation and translation, from one point cloud (source) to the other (target). It has attracted a lot of research interest due to its broad applications in SLAM (Simultaneous Localization and Mapping) [16, 50], autonomous driving [34, 37], 3D reconstruction [26, 22], etc.
Figure 1: Our approach focuses on directly matching the predicted superpoints by computing correlation between them. The bold dots show the superpoints with highest correlation which are used as correspondences to correctly register the point cloud pair.
A prevailing paradigm to solve the registration task is to leverage the correspondences of superpoints across two point clouds, which can be obtained either using separate keypoint detectors [1, 4, 13, 24] to capture salient and distinctive points or regions within a point cloud. With the rapid advancement in representation learning in point clouds using deep neural networks [42, 43, 51], the keypoint-free approaches have gained significant attention, where the point cloud is downsampled into superpoints. Although they may not be designed to capture the salient keypoints in the point cloud, the learned representations endow the superpoints with discriminative power so they can be matched across the source and target point clouds. The transformation matrix can then be obtained based on the superpoints correspondences using the Kabsh-Umeyama Algorithm [53, 28].
Direct superpoint matching is typically not used by itself in existing approaches for point cloud registration for two reasons. First, the correspondences of superpoints inevitably contain errors, either because the point clouds contain noisy sensory data or because some of the correspondences are simply incorrect. As a result, postprocessing refinement is usually needed, for instance, using RANSAC to prune out the outliers [13, 14, 68]. However, RANSAC is inherently slow and not differentiable, and thus cannot be integrated into the training step. Second, superpoints are essentially sparse, particularly because of downsampling in neural networks. Therefore, more RANSAC iterations are needed as the number of inliers (i.e., correct correspondences) is small, which imposes a significant computation burden. To this end, the coarse-to-fine registration scheme is usually adopted in state-of-the-art approaches [44, 66, 38], where superpoints correspondences serve as coarse correspondences only. They are then propagated to the point level, forming finer-level correspondences. Since point-wise features only capture local information, they lack enough discriminative power. Consequently, the noise in the point-wise correspondences is more prominent, which further necessitates the additional refinement step. Although the Local to Global Registration (LGR) module introduced in [44] improves the speed, which works similarly to RANSAC in an iterative manner, the selection and refinement of the transformation still remain non-differentiable. The entire model thus can not be optimized jointly, leading to inferior accuracy.
In this paper, we present a simple yet effective approach for rigid point cloud registration by directly matching the superpoints. Specifically, superpoints features are first obtained from the KPConv backbone [51], which are then enhanced using Transformer blocks with interleaved multi-head self and cross attention modules [64], enabling effective learning of the structural information of from the target to the source point cloud. The enhanced superpoint features are finally matched using a Global Softmax layer inspired by the 2D matching work [60], which generates a correlation matrix capturing the pairwise relationships between the keypoints from the source and target point clouds. But unlike [60], we fully harness the rich information in the correlation matrix. We employ a differentiable variant of the Kabsh-Umeyama Algorithm [53, 28], considering the strengths of matchings, where the weights of the SVD (Singular Value Decomposition) are the correlation scores between two corresponding superpoints. By examining the correlation scores of the superpoints, we can easily identify the highly reliable correspondences without resorting to the time-consuming RANSAC or the non-differentiable LGR module [44]. Our approach is fully differentiable and can be trained in an end-to-end manner, which enables joint optimization of the feature representation learning, superpoints matching, and transformation estimation, leading to improved alignment results. Moreover, our approach does not require ad-hoc postprocessing, e.g., RANSAC, it runs fast, making it potentially more useful in practice. Compared with the recent end-to-end method [64], our approach directly matches the superpoints. With the rich information obtained from the matching step, our approach can get a more accurate estimation of the transformation between the point clouds robustly by filtering out the outliers effectively.
We run experiments on standard benchmarks, including ModelNet [58] and 3DMatch [68], and achieve state-of-the-art results. Extensive ablation studies validate the effectiveness of each module in our proposed approach. The key contributions of the paper can be summarized as follows:
  • Efficiency. Our approach tackles the point cloud registration by directly matching superpoints. The rich information in the matching step allows us to effectively filter out outliers without using RANSAC-like approaches. As a result, our approach is efficient and runs fast in practice.
  • Simplicity. Our registration pipeline can be trained end-to-end without the any ad hoc postprocessing refinement. It streamlines the process and makes it more straightforward to implement and understand.
  • Robustness. We thoroughly evaluate our method on various datasets, demonstrating its superior performance in comparison to existing techniques. Our approach achieves state-of-the-art results on multiple benchmarks, with notably high inlier ratios. This showcases the robustness and accuracy of our proposed registration framework.
2 Related Work
Traditional registration approaches. The most known algorithm Iterative Closest Point (ICP) [5] has been widely used for point cloud registration. ICP solves the registration problem iteratively in two steps: (1) It obtains the spatially closest point correspondence and then (2) finds the least-squares rigid transformation. The spatial-distance-based correspondences are sensitive to the initial transformation and point noises. A lot of variants [46, 11, 49, 6, 45, 35] have been proposed to improve ICP. Another line of work [27, 17] formulates the point cloud registration as probability distribution matching problems to improve the robustness and to converge quickly. However, they still heavily rely on an appropriate initialization and can easily converge to a local optimum as opposed to the global solution. Unlike the local methods mentioned above, global approaches are invariant to the initialization. Based on branch-and-bound techniques, several works [61, 8, 9, 35] have been proposed, but they are very slow and impractical in some scenarios. Another method is to extract and match keypoints based on feature extraction methods such as FPFH [47] and SHOT [52]. Then RANSAC [18] can be used for registration. However, RANSAC is computationally very slow compared to ICP.
Learning-based registration approaches. Recently, many works have used deep learning for point cloud learning and registration. Some work first estimates the correspondence between two point clouds and then computes the transformation with some robust pose estimators. To predict the correspondence between two point clouds, 3DMatch [68] detects the repeatable keypoints and learns discriminative descriptors for keypoints. The following works aim to either improve the keypoint detections [4, 32, 62] or learn better feature descriptors [13, 14, 15, 29, 1, 55]. Predator [24] uses the attention mechanism proposed in Transformers [54] to enhance the point feature descriptors. Other detector-free methods [66, 44] extract the correspondences by considering all possible matches. Another line of work [10, 67] has included the transformation computation into the training pipeline. Following the idea of ICP [5],  [57, 56, 63, 19, 31] iteratively predict the soft correspondences and computes the transformation with differentiable weighted SVD. Another way [2, 33, 25, 59, 48] first extracts a global feature vector for each point cloud and predicts the transformation with global feature vectors. This approach usually fails in large-scale scenes. Unlike these works which require eighter ad-hoc postprocessing or coarse-to-fine registration, our method directly matches the superpoints without any refinement.
Correspondence filters. RANSAC [18] is typically used to filter out the outliers in the predicted correspondence to obtain a robust transformation estimation. However, RANSAC is relatively slow and cannot be incorporated into the training pipeline because the hypothesis selection step is non-differentiable. To alleviate these problems, DSAC [7] modify the RANSAC pipeline and make it differentiable. Other deep robust estimators [3, 12, 40, 20, 30, 65] usually use the classification network to identify which correspondences are outliers and then reject them. Instead of using these complex correspondence filters, our method can directly filter out outliers effectively by leveraging the rich information in the superpoints matching.
3 Method
Given the source and target point clouds $\mathbf{X}\in\mathbb{R}^{M\times 3}$ and $\mathbf{Y}\in\mathbb{R}^{N\times 3}$, our goal is to determine the optimal rigid transformation $\mathbf{T}=\{\mathbf{R},\mathbf{t}\}$ with rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$ to align two point clouds into a common coordinate system. $M$ and $N$ denote the numbers points.
3.1 Superpoints Feature Extraction and Enhancement
In our approach, we use Kernel Point Convolution (KPConv) [51] as the backbone to selectively downsample the point cloud into a set of superpoints and extract global feature vectors for each superpoint. Specifically, the KPConv backbone uses a series of ResNet-like blocks [23] and convolutions to downsample the input point clouds into a reduced set of superpoints $\mathbf{X}{{}^{\prime}}\in\mathbb{R}^{M^{\prime}\times 3}$ and $\mathbf{Y}{{}^{\prime}}\in\mathbb{R}^{N^{\prime}\times 3}$, where $M^{\prime}<M$ and $N^{\prime}<N$. The superpoints are described by their feature vectors $\mathbf{F}_{\mathbf{X}{{}^{\prime}}}\in\mathbb{R}^{M^{\prime}\times D}$ and $\mathbf{F}_{\mathbf{Y}{{}^{\prime}}}\in\mathbb{R}^{N^{\prime}\times D}$, respectively, with $D$ being the feature dimension. The network weights are shared among the two point clouds. We use a shallower backbone for 3DMatch dataset compared to [64, 24] to avoid significant downsampling.
@@ 3.1 Superpoints Feature Extraction and Enhancement
Figure 2: Model Architecture: The KPConv backbone downsamples the input point cloud and generates superpoints and superpoint feature vectors. These superpoint features are then conditioned on the other point cloud in the Feature Enhancement block. Lastly, superpoint features can be directly matched using Global Softmax to predict correspondences and estimate rigid transformation.
@@
Although KPConv backbone provides reasonably good representations, the superpoints features are obtained within each point cloud independently. To obtain highly discriminative feature representations for superpoints matching, we further enhance their feature representations in the source point cloud to be conditioned on the target one and vice-versa. Following the previous work of point cloud registration [64] and the 2D counterpart of superpoints matching (e.g., optical flow) [60], we adopt the multi-head attention mechanism in the Transformer model [54] as the feature enhancement module, shown in Fig. 2. It consists of both self and cross-attention, where the self-attention is to integrate the information from the other points within the same point cloud and the cross-attention allows interactions with points in another point cloud to consider the mutual dependencies. In addition to the multi-head attention, other components in the Transformer model, including position encodings of 3D points, residual connections, layer normalization, and feed-forward network are applied to each layer. The entire feature enhancement module consists of 6 such layers with 256 dimensions and 8 attention heads.
@@
The outputs of the feature enhancement module are features $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}\in\mathbb{R}^{M^{\prime}\times D}$ and $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}\in\mathbb{R}^{N^{\prime}\times D}$ which has aggregated geometric information from both source and target point cloud. The strongly associated features are strengthened while the weakly associated features are weakened. With such highly discriminative feature representations, we can obtain high-quality superpoints matchings.
3.2 Direct Superpoint Matching for Rigid Transformation Estimation
@@
To get the correspondences between two point clouds, we first compare the feature similarity for each point in $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ to all points in $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$ by computing their correlations [60], which can be done efficiently in a single step as follows:
@@
${\mathbf{C}}=\mathrm{softmax}\left(\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}^{T}\right)\in\mathbb{R}^{M^{\prime}\times N^{\prime}},$ (1)
@@
where $\mathbf{C}$ is the normalized correlation matrix that represents the similarity between two point clouds. Based on the correlation matrix, the correspondences $\mathbf{\hat{Y}}$ and $\mathbf{\hat{X}}$ can be directly calculated by using the largest correlation for each point.
@@
The rigid transformation between the source and target point clouds can be then estimated using the superpoints correspondences with a weighted variant of the Kabsch-Umeyama algorithm [28, 53]:
@@
where $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{y}}_{i}$ are two matched superpoints, denoting the $i^{th}$ point in $\mathbf{\hat{X}}$, $\mathbf{\hat{Y}}$, respectively. The coefficient $w_{i}$ can be used to weigh different correspondences.
Although we use the highly discriminative feature representations enhanced by the attention module, the correspondences are inevitably noisy. How to filter out the outliers (i.e., incorrect correspondences)? We show that the normalized correlation matrix $\mathbf{C}$ obtained from superpoints matching in Eq. (1) contains rich information, which allows us effectively select high-confident superpoints correspondences to estimate the transformation. Specifically, if $\hat{\mathbf{x}}_{i}$ is similar to multiple superpoints, e.g., because of the repetitive patterns, its matching to $\hat{\mathbf{y}}_{i}$ tends to be unreliable. Therefore, the normalized correlation score between them will be low since $\mathbf{C}$ is normalized w.r.t. all other superpoints in the target point cloud. We can therefore set $w_{i}=\mathbf{C}(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})$, capturing the confidence of the superpoints matchings. On the other hand, we can use $w_{i}$ to select only highly confident superpoint matchings, e.g., top 15% of them, to estimate the transformation in Eq.(2. We can also augment the weights $w_{i}$ by considering the overlap score, which is introduced below, to further improve the accuracy.
Compared to previous work [44, 66, 38], our approach is simple yet effective, which does not only eliminates the coarse-to-fine strategy but more importantly, the slow and non-differentiable RANSAC. As a result, our approach is fast and the superpoints feature learning, superpoints matching, and transformation estimation can be optimized jointly in an end-to-end manner, leading to superior accuracy. Although the feature enhancement module is also used in [64], our approach is fundamentally different. Instead of predicting the correspondences by using the feature representations of the source point cloud only, we perform superpoints matching and use the rich information in the matching step to effectively filter outliers and get a robust estimation of the transformation.
3.3 Loss Functions

### v4

# Direct Superpoints Matching for Robust Point Cloud Registration

@@ Abstract
Deep neural networks endow the downsampled superpoints with highly discriminative feature representations. Previous dominant point cloud registration approaches match these feature representations as the first step, e.g., using the Sinkhorn algorithm. A RANSAC-like method is then usually adopted as a post-processing refinement to filter the outliers. Other dominant method is to directly predict the superpoint matchings using learned MLP layers. Both of them have drawbacks: RANSAC-based methods are computationally intensive and prediction-based methods suffer from outputing non-existing points in the point cloud. In this paper, we propose a straightforward and effective baseline to find correspondences of superpoints in a global matching manner. We employ the normalized matching scores as weights for each correspondence, allowing us to reject the outliers and further weigh the rest inliers when fitting the transformation matrix without relying on the cumbersome RANSAC. Moreover, the entire model can be trained in an end-to-end fashion, leading to better accuracy. Our simple yet effective baseline shows comparable or even better results than state-of-the-art methods on three datasets including ModelNet, 3DMatch, and KITTI. We do not advocate our approach to be the solution for point cloud registration but use the results to emphasize the role of matching strategy for point cloud registration. The code and models are available at https://github.com/neu-vi/Superpoints_Registration
I Introduction
@@
Point cloud registration refers to the task of aligning two partially overlapping point clouds into a shared coordinate system. In this paper, we tackle this problem where the goal is to determine the transformation matrix, including rotation and translation, from one point cloud (source) to the other (target). It has attracted a lot of research interest due to its broad applications in SLAM (Simultaneous Localization and Mapping) [14, 44], autonomous driving [28, 32], 3D reconstruction [22, 19], etc.
Figure 1: Our approach directly matches the superpoints between the input point clouds to register them by estimating the $SE(3)$ transformation matrix. The correlation weights obtained from matching are used to filter out incorrect correspondences (outliers) and further weigh the rest inliers for the transformation matrix estimation. The source cloud is shown in blue, with superpoints highlighted in red. Similarly, the target point cloud is displayed in yellow, with matching superpoints shown in blue. The intensity of the red and blue colors represents the correspondence weights. Notice how most keypoints are distributed around the three-way junction and exhibit high correspondence weights. (Best viewed in color.)
A prevailing paradigm to solve the registration task is to leverage the correspondences of superpoints across two point clouds, which can be obtained using either keypoint detectors [1, 4, 11, 21] or downsampling in deep neural networks [37, 38, 47]. Ideally, these superpoints should capture salient and distinctive points or regions within a point cloud. With the learned feature representations, the superpoints are endowed with sufficient discriminative power so they can be matched across the source and target point clouds, where the matching strategy is crucial for finding correspondences of superpoints and can significantly affect the convergence and performance of the network. Previous approaches  [58, 39] use the Sinkhorn algorithm [45] to match superpoints, which however is sensitive to initialization parameters and requires careful tuning. In  [46], Dual-Softmax is adopted for feature matching between two input images by using kepoints which have softmax correlation scores greater than a certain threshold. The correspondences of superpoints, however, are inevitably noisy due to either the noisy sensory data or simply incorrect estimations of correspondences, which prevent existing approaches from directly matching superpoints for point cloud registration. Postprocessing refinement is usually needed, for instance, using RANSAC-like approaches to prune the outliers [11, 12, 63, 61]. But such refinement is inherently slow due to the iterative nature of the RANSAC pipeline and can not be easily integrated into an end-to-end trainable system.
The recent work [59] populates another strategy, which uses a two-layer MLP to direct predict matched superpoints between the input point clouds. But such predicted correspondences may not exist in the point cloud, leading to inferior registration accuracy.
In this paper, we present a strong baseline for point cloud registration by directly matching superpoints. Specifically, building upon RegTR [59], instead of using a MLP to predict correspondences of superpoints, we find the correspondences by computing the similarity scores of all superpoints across the source and target point cloud in a global matching manner. Their normlized matching scores can be used to filter out the unreliable correspondences (i.e., outliers). By integrating the weights of the rest inlier superpoint matchings into a differentiable variant of the Kabsh-Umeyama Algorithm [50, 23], we obtain robust estimations of the $SE(3)$ transformation between the input point cloud pairs. An illustration is shown in Fig. 1. As a result, no ad hoc postprocessing refinement is needed, yielding a more efficient model. More importantly, the entire model can be trained in an end-to-end manner, where the feature representation learning, superpoints matching, and transformation estimation can be jointly optimized. Better registration accuracy can thus be obtained. In addition, compared to [59], our approach find correspondences of superpoints by global matching, which does not output non-existing points in the target point cloud.
We run experiments on three benchmark datasets, including ModelNet [54], 3DMatch [63], and KITTI [16], and achieve comparable or even better results than state-of-the-art methods. We do not advocate our approach to be the solution for point cloud registration. Rather, we’d like to emphasize the role of matching strategy for point cloud registration, showcasing the feasibility of achieving high accuracy without cumbersome ad hoc postprocessing. By releasing the code and model weights, we hope our work can foster future research.
@@
Figure 2: Model Architecture: The KPConv backbone downsamples the input point cloud and generates superpoints and their feature vectors. These superpoint features are then conditioned on the other point cloud in the feature enhancement block. Lastly, superpoint features can be directly matched using Global Softmax to estimate SE(3) transformation using the corrleation weights in a robust manner. The intensity of the red and blue colors of the superpoints after the softmax correlation step represents the correspondence weights. (Best viewed in color.)
II Related Work
Traditional registration approaches. The most known algorithm Iterative Closest Point (ICP) [5] has been widely used for point cloud registration. ICP solves the registration problem iteratively in two steps: (1) It obtains the spatially closest point correspondence and then (2) finds the least-squares rigid transformation. The spatial-distance-based correspondences are sensitive to the initial transformation and point noises. A lot of variants [41, 9, 43, 6, 40, 30] have been proposed to improve ICP. Another method is to extract and match keypoints based on feature extraction methods such as FPFH [42] and SHOT [48], followed by an outlier rejection postprocessing step.
Learning-based registration approaches. Recently, many works have used deep learning for point cloud learning and registration. Some work first estimates the correspondence between two point clouds and then computes the transformation with some robust pose estimators. To predict the correspondence between two point clouds, 3DMatch [63] detects the repeatable keypoints and learns discriminative descriptors for keypoints. The following works aim to either improve the keypoint detections [4, 26, 57] or learn better feature descriptors [11, 12, 13, 24, 1, 52]. Predator [21] uses the attention mechanism proposed in Transformers [51] to enhance the point feature descriptors. Other detector-free methods [61, 39] extract the correspondences by considering all possible matches. Another line of work [8, 62] has included the transformation computation into the training pipeline. Unlike these works which require either ad-hoc postprocessing or coarse-to-fine registration, our method directly matches the superpoints without any refinement.
Learning visual correspondence. Various approaches have been proposed to establish correspondences between the input, e.g., using nearest neighbor followed by the distance ratio test [29]. Recent approaches [58, 39] use the Sinkhorn algorithm [45] to match superpoints for point cloud registration. However, it is sensitive to initialization parameters and requires careful tuning. In  [46], Dual-Softmax is adopted for feature matching between two input images by using kepoints which have softmax correlation scores greater than a certain threshold. Global Softmax is used in a recent work of optical flow [56] to find global matchings by simply taking the correspondence for each pixel with the highest correlation score. In this paper, we extend Global Softmax by using the corrleation weights for both filtering unreliable correspondences and weighing the remaining inliers when estimating the SE(3) transformation matrix.
Correspondence filters. RANSAC [15] is typically used to filter out the outliers in the predicted correspondence to obtain a robust transformation estimation. However, RANSAC is relatively slow and cannot be incorporated into the training pipeline because the hypothesis selection step is non-differentiable. To alleviate these problems, DSAC [7] and $\nabla$-RANSAC [53] modify the RANSAC pipeline and make it differentiable. But even the differentiable versions are similar in computational complexity as vanilla RANSAC. Other deep robust estimators [3, 10, 35, 17, 25, 60] usually use the classification network to identify which correspondences are outliers and then reject them. Instead of using these complex correspondence filters, our method can directly filter out outliers effectively by leveraging the rich information in the superpoints matching.
III Method
Given the source and target point clouds $\mathbf{X}\in\mathbb{R}^{M\times 3}$ and $\mathbf{Y}\in\mathbb{R}^{N\times 3}$, our goal is to determine the SE(3) transformation $\mathbf{T}=\{\mathbf{R},\mathbf{t}\}$ with rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$ to align two point clouds into a common coordinate system. $M$ and $N$ denote the numbers of points.
III-A Superpoints Feature Extraction and Enhancement
Following [59], we use Kernel Point Convolution (KPConv) [47] as the backbone to selectively downsample the point cloud into a set of superpoints and extract global feature vectors for each superpoint. The KPConv backbone uses a series of ResNet-like blocks [20] and convolutions to downsample the input point clouds into a reduced set of superpoints $\mathbf{X}{{}^{\prime}}\in\mathbb{R}^{M^{\prime}\times 3}$ and $\mathbf{Y}{{}^{\prime}}\in\mathbb{R}^{N^{\prime}\times 3}$, where $M^{\prime}<M$ and $N^{\prime}<N$. The superpoints are described by their feature vectors $\mathbf{F}_{\mathbf{X}{{}^{\prime}}}\in\mathbb{R}^{M^{\prime}\times D}$ and $\mathbf{F}_{\mathbf{Y}{{}^{\prime}}}\in\mathbb{R}^{N^{\prime}\times D}$, respectively, with $D$ being the feature dimension. The network weights are shared among the two point clouds. We use a shallower backbone for 3DMatch dataset compared to [59, 21] to avoid significant downsampling by removing the 4-th residual block.
@@ III-A Superpoints Feature Extraction and Enhancement
Although KPConv backbone provides reasonably good representations, the superpoints features are obtained within each point cloud independently. To obtain highly discriminative feature representations for superpoints matching, We use the multi-head attention mechanism in the Transformer model [51] as the feature enhancement module as suggested in [59], shown in Fig. 2. It consists of both self and cross-attention, where the self-attention is to integrate the information from the other points within the same point cloud and the cross-attention allows interactions with points in another point cloud to consider the mutual dependencies. In addition to the multi-head attention, other components in the Transformer model, including position encodings of 3D points, residual connections, layer normalization, and feed-forward network are applied to each layer. The entire feature enhancement module consists of 6 such layers with 256 dimensions and 8 attention heads.
@@
The outputs of the feature enhancement module are features $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}\in\mathbb{R}^{M^{\prime}\times D}$ and $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}\in\mathbb{R}^{N^{\prime}\times D}$ which has aggregated geometric information from both source and target point cloud. The strongly associated features are strengthened while the weakly associated features are weakened.
III-B Superpoint Matching for SE(3) Transformation Estimation
@@
Unlike [59] that uses a two-layer MLP to predict corresponding points, we find the correspondences of superpoints in a global matching manner. This change removes the ambiguity of predicting non-existing corresponding points. To get the correspondences between two point clouds, we first compare the feature similarity for each point in $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ to all points in $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$ by computing their correlations [56], which can be done efficiently in a single step as follows:
@@
${\mathbf{C}}=\mathrm{Softmax}\left(\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}\cdot\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}^{T}\right)\in\mathbb{R}^{M^{\prime}\times N^{\prime}},$ (1)
@@
where $\mathbf{C}$ is the normalized correlation matrix representing the similarity between two point clouds. Based on the correlation matrix, the correspondences between $\mathbf{X}{{}^{\prime}}$ and $\mathbf{Y}{{}^{\prime}}$ can be directly calculated by using the largest correlation for each point.
@@
The SE(3) transformation between the source and target point clouds can be then estimated using the superpoints correspondences with a weighted variant of the Kabsch-Umeyama algorithm [23, 50]:
@@
where $w_{i}=\mathbf{C}(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})$ is the correspondence weight, $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{y}}_{i}$ are the $i$-th pair of matched superpoints. We use $\min(N^{\prime},M^{\prime})$ operation because corresponding points can only be found for the point cloud with the minimum number of superpoints.
Although we use the highly discriminative feature representations enhanced by the attention module, the correspondences are inevitably noisy. How to filter out the outliers (i.e., incorrect correspondences)? We show that the normalized correlation matrix $\mathbf{C}$ obtained from superpoints matching in Eq. (1) contains rich information, allowing us to effectively reject outliers for robust transformation matrix estimation. Specifically, if $\hat{\mathbf{x}}_{i}$ is similar to multiple superpoints, e.g., because of the repetitive patterns, its matching to $\hat{\mathbf{y}}_{i}$ tends to be unreliable. Therefore, the normalized correlation score between them $w_{i}=\mathbf{C}(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})$ will be low since $\mathbf{C}$ is normalized w.r.t. all other superpoints in the target point cloud. Unlike [56], which completely discards the correlation weights in $\mathbf{C}$, we use $w_{i}$ for two purposes: discarding the unreliable corresondences that have low $w_{i}$ values and further weighting the rest inliers when fitting the transformation matrix in Eq. (1). It is important to note here that $w_{i}$ is not learned and is just the correlation score between the feature representations of $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{y}}_{i}$, the $i$-th pair of matched superpoints. We show in the experiments that such a matching strategy works more effectively for superpoints matching than other approaches [49, 56, 46].
In comparison to  [39, 61, 33], our approach is simple yet effective, which eliminates the coarse-to-fine strategy and more importantly, the inherently slow RANSAC-like postprocessing. Although the feature enhancement module is also used in [59], our approach is fundamentally different.  [59] predicts corresponding points for both source and target point clouds whereas our approach removes the ambiguity of predicting non-existing points and directly matches superpoints. The limitation of [59] can also be seen in outdoor LiDAR experiments like KITTI. As shown in Table III, it performs significantly worse due to predicting correspondences that do not align well with true corresponding points.
III-C Loss Functions